# Diffusion-Based Impedance Learning for Contact-Rich Manipulation Tasks

### v2

@@ III-D3 Algorithm
1:Input: $\begin{aligned} &\bm{e}_{t},\bm{e}_{r},\dot{\bm{e}},\bm{\omega},\bm{f}_{\text{ext}},\bm{m}_{\text{ext}},\kappa_{t},\kappa_{r},\gamma_{t},\gamma_{r},\\ &f_{\text{thres}},m_{\text{thres}},\bm{K}_{t,\text{max}},\bm{K}_{r,\text{max}},\varepsilon\end{aligned}$
@@
3:  $\tilde{e}_{t,i}\leftarrow\kappa_{t}\,e_{t,i}-\gamma_{t}\,\dot{e}_{i}$
@@
11:  $\tilde{e}_{r,i}\leftarrow\kappa_{r}\,e_{r,i}-\gamma_{r}\,\omega_{i}$
@@
III-D4 Directional adaptation
For deployment of our model in real world experiments, the energy-based estimator was extended with a directional adaptation step. A uniform reduction of stiffness in response to external perturbations can prevent task execution, e.g., the end-effector may stop when encountering an obstacle. To avoid this problem, stiffness was modulated according to the directional relevance of each axis with respect to the model-generated sZFT. The alignment factors below are defined relative to $\bm{p}_{0}$ and $\bm{u}_{0}\theta_{0}$ contained in $\bm{e}_{t}$ and $\bm{e}_{r}$, which are outputs of the Diffusion Model. Hence, the adaptation mechanism depended critically on the reconstructed sZFT: if the nominal free-space ZFT were used directly, the alignment factors would not reflect the equilibrium under contact, and directional adaptation would fail (Section IV).
@@ III-D4 Directional adaptation
$\psi_{t,i}=\frac{|e_{t,i}|}{\lVert\bm{e}_{t}\rVert},\quad i\in\{x,y,z\}$ (25)
quantifies the contribution of each axis to the intended displacement. Analogously, for rotations we define
$\psi_{r,i}=\frac{|e_{r,i}|}{\lVert\bm{e}_{r}\rVert},\quad i\in\{x,y,z\}.$ (26)
@@
The final adapted stiffness values $k_{t,i}$ and $k_{r,i}$ were implemented as
@@ IV Deployment of Diffusion Model on Real Robot
The model was deployed in two robotic use-cases designed to evaluate its performance in contact-rich scenarios. We deliberately focus on Impedance Control rather than motion generation, since many real world applications require active compliance, e.g., final assembly tasks where vision is occluded and only tactile sensing is available. The first was a parkour-style task requiring continuous interaction while traversing unknown obstacles. The second was a peg-in-hole insertion task with increasing geometric complexity.
The experiments were conducted on a KUKA LBR iiwa with seven DOFs, using KUKA’s Fast Robot Interface (FRI) for torque control with 5 ms sample time. Built-in gravity and Coriolis/centrifugal compensation remained active throughout. Kinematic and dynamic matrices were computed via the Exp[licit]-FRI interface [^7] https://GitHub.com/explicit-robotics/Explicit-FRI [58]. The external wrenches were aquired with an ATI Gamma force/torque transducer, mounted on the robot flange. The code is publicly available on our GitHub repository [^8] https://github.com/StrokeAIRobotics/DiffusionBasedImpedanceLearning.
@@
All workpieces were fabricated with a PRUSA i3 MK3 3D printer using PLA filament, and CAD models are available on our GitHub repository. Demonstration videos of both experiments are provided on the project website [^9] https://strokeairobotics.github.io/DiffusionBasedImpedanceLearning.
@@ IV-A Parkour
The first experiment was a parkour-style task in which the robot had to traverse multiple obstacles with continuous contact. The nominal parkour ZFT was generated from a sequence of key configurations connected via minimum-jerk and SLERP interpolation (see code on GitHub).
@@
To protect the setup, stop conditions of $\lVert\bm{v}\rVert=0.24\ \frac{\text{m}}{\text{s}}$ and $\lVert\bm{f}_{\text{ext}}\rVert=20$ N were defined. The constant-stiffness controller already failed at the first obstacle: the end-effector was stopped by the bumper, energy accumulated in the virtual spring, and once this energy was released the end-effector accelerated into the velocity stop condition. Forces of up to 15 N were observed. In contrast, with diffusion-based impedance adaptation, the robot traversed all obstacles smoothly, i.e., not violating the velocity or force stop conditions. The stiffness adaptation mechanism prevented jamming in directions inconsistent with the ZFT while preserving rigidity along task-relevant axes.
The videos of the experiments with constant sitffness and Diffusion-Based Impedance Learning can be found on our project website.
@@
Figure 8 shows the measured external wrenches together with the commanded translational and rotational stiffness values for the Diffusion-Based Impedance Learning. The bottom plots illustrate the relative translational and rotational factors (Equation (27)) that modulate the stiffness based on the contribution of each axis to the ZFT.
@@
Figure 8: Measured external wrenches and commanded translational and rotational stiffness during the parkour experiment with stiffness adaptation. All quantities are expressed in the robot base frame. The lower plots show the relative translational and rotational factors and their influence on stiffness modulation. The highlighted gray regions illustrate stiffness reduction in response to external wrenches, governed by the relative translational and rotational factors.
@@
During the entire motion, stiffness values changed dynamically in response to external wrenches. This continuous adaptation reflected the underlying Diffusion-Based Impedance Learning, which enabled the robot to remain compliant when unexpected contact occurred while maintaining sufficient stiffness to progress through the parkour.
@@
Translational stiffness adapted asymmetrically depending on the directional contribution of forces. Around 12–13 s, similar forces appeared along both the $y$- and $z$-axes. Nevertheless, $K_{p,z}$ was reduced more strongly (to about 200 N/m) than $K_{p,y}$ (remaining near 500 N/m). A similar effect is visible near 30 s: while $f_{y}$ reached about 4 N with stiffness near 400 N/m, $f_{z}$ was only 2 N yet stiffness dropped to about 200 N/m. These findings confirm that stiffness reduction was not simply proportional to force magnitude but was determined by the relative translational factors and their alignment with the sZFT.
@@
Rotational stiffness also adapted directionally, with reductions depending on which axis contributed less to the sZFT. Around 19 s, moments appeared simultaneously about the $x$- and $y$-axes. Despite higher magnitudes about the $x$-axis, the stiffness $K_{r,y}$ was reduced more strongly, showing that $y$-axis rotations were less important for the intended trajectory. This selective modulation allowed the robot to maintain stability while avoiding unnecessary resistance to contact.
@@ IV-B Peg-In-Hole Insertion
The second experiment was a peg-in-hole insertion task with three different peg types: cylindrical, square, and star pegs. The nominal clearance between each peg and its corresponding hole was 0.14 mm for the square peg and 0.20 mm for the circular and star pegs. All workpieces were fabricated with the same printer setup as for the parkour experiment. For all trials, task success was defined as completely inserting the peg.
As with the parkour experiment, nominal trajectories were constructed by minimum-jerk interpolation between 3 key robot poses: approach and contact, final insertion, and retreat.
@@
The three peg types represent increasing geometric complexity. The cylindrical peg is rotationally symmetric, which makes the final orientation about the end-effector $z$-axis irrelevant. In contrast, the square and star pegs require precise alignment with an increasing number of edges, which significantly increases task complexity. This was confirmed by the experiments with constant stiffness values: the cylindrical peg was inserted successfully in 30/30 trials, the square peg in 4/30 trials, and the star peg in none of the 30 trials.
In contrast, Diffusion-Based Impedance Learning achieved a 30/30 success rate for all three peg types. These results highlight that with increasing insertion complexity, more advanced strategies such as our diffusion-based approach become essential. The outcome is particularly notable because the training data included only parkour and upper-limb rehabilitation data (Section III-A) and no peg-in-hole demonstrations.
@@
The videos of the experiments with constant sitffness and Diffusion-Based Impedance Learning can be found on our project website [^10] https://strokeairobotics.github.io/DiffusionBasedImpedanceLearning.
@@
Figure 9 shows the stiffness adaptation for the star peg. Results for the other two pegs are provided in Appendix C. During all experiments, force and moment thresholds were set to 11 N and 11 Nm, respectively.
@@
Figure 9: Measured external wrenches and commanded translational and rotational stiffness during the star peg insertion experiment. All quantities are expressed in the robot base frame. The lower plots show the directional motion factors for translation and rotation and their influence on stiffness modulation. The highlighted gray region illustrates stiffness reduction in response to external wrenches, governed by the relative translational and rotational factors.
@@
As illustrated in Figure 9, stiffness values were reduced in response to external wrenches, with the amount of reduction governed by the relative translational and rotational factors. At around 11 s, forces of 3–4 N appeared along both the $y$- and $z$-axes. Nevertheless, the stiffness $K_{p,z}$ was reduced more strongly than $K_{p,y}$ because the relative translational factor in $z$ was lower. At around 13 s, even though a force peak of 14 N occurred along $z$, the stiffness was only marginally reduced since the relative translational factor approached 1.
Due to the high stiffness reduction of translations, the external moments remained small (below 4 Nm). The effect of the relative rotational factor can be seen for rotations about the $x$- and $y$-axes near 12 s. A moment of 4 Nm about the $x$-axis and 1 Nm about the $y$-axis resulted in a stronger reduction of $K_{r,y}$ than $K_{r,x}$, reflecting the higher contribution of $y$-axis rotations to the sZFT.
@@
Ablation study. In addition to the two main controllers, we evaluated uniform stiffness adaptation and directional adaptation computed from the nominal ZFT (i.e., the commanded free-space trajectory). Uniform stiffness adaptation failed in both scenarios: the parkour traversal was blocked at the first obstacles, and none of the peg-in-hole insertions succeeded. Directional adaptation from the nominal ZFT also failed in contact-rich settings: in parkour, the robot hit the safety bounds and stopped, while in peg-in-hole tasks it achieved 30/30 success only for the cylindrical peg but failed for both square and star pegs. These results confirm that directional factors must be derived from the reconstructed, contact-consistent sZFT rather than from the nominal ZFT in order for impedance adaptation to succeed.

### v3

@@ III-E3 Algorithm
1:Input: $\begin{aligned} &\hat{\bm{e}}_{t},\hat{\bm{e}}_{r},\dot{\bm{e}},\bm{\omega},\bm{f}_{\text{ext}},\bm{m}_{\text{ext}},\kappa_{t},\kappa_{r},\gamma_{t},\gamma_{r},\\ &f_{\text{thres}},m_{\text{thres}},\bm{K}_{t,\text{max}},\bm{K}_{r,\text{max}},\varepsilon\end{aligned}$
@@
3:  $\tilde{e}_{t,i}\leftarrow\kappa_{t}\,\hat{e}_{t,i}-\gamma_{t}\,\dot{e}_{i}$
@@
11:  $\tilde{e}_{r,i}\leftarrow\kappa_{r}\,\hat{e}_{r,i}-\gamma_{r}\,\omega_{i}$
@@
III-E4 Directional Adaptation
For deployment of our model in real-world experiments, the energy-based estimator was extended with a directional adaptation step. A uniform reduction of stiffness in response to external perturbations can prevent task execution, e. g., the end-effector may stop when encountering an obstacle. To avoid this problem, stiffness was modulated according to the directional relevance of each axis with respect to the generated sZFT of the Diffusion Model. The alignment factors below are defined relative to $\hat{\bm{p}}_{0}$ and $\hat{\bm{u}}_{0}\hat{\theta}_{0}$ contained in $\hat{\bm{e}}_{t}$ and $\hat{\bm{e}}_{r}$, which are outputs of the Diffusion Model. Hence, the adaptation mechanism depended critically on the reconstructed sZFT: if the nominal free-space ZFT were used directly, the alignment factors would not reflect the equilibrium under contact, and directional adaptation would fail (Section IV).
@@ III-E4 Directional Adaptation
$\psi_{t,i}=\frac{|\hat{e}_{t,i}|}{\lVert\hat{\bm{e}}_{t}\rVert},\quad i\in\{x,y,z\}$ (32)
quantifies the contribution of each axis to the intended displacement. Analogously, for rotations, we define
$\psi_{r,i}=\frac{|\hat{e}_{r,i}|}{\lVert\hat{\bm{e}}_{r}\rVert},\quad i\in\{x,y,z\}.$ (33)
@@
The final adapted stiffness values $k_{t,i}$ and $k_{r,i}$ were implemented as:
@@
An overview of the proposed model architecture and workflow can be seen in Figure 8.
Figure 8: Overview of the proposed architecture and workflow. External wrenches and end-effector pose (position and unit-quaternion) are processed as time-series inputs to a conditional Transformer-based Diffusion Model. External wrenches are embedded as context tokens, while pose tokens are embedded with learned positional and diffusion-time encodings. The Diffusion Model reconstructs a simulated Zero-Force Trajectory (sZFT), representing a contact-consistent equilibrium. This reconstructed sZFT, together with the measured interaction signals, is passed to an energy-based stiffness estimator. Finally, directional stiffness adaptation is applied according to the directional relevance of each coordinate axis.
@@ IV Deployment of Diffusion Model on Real Robot
The model was deployed in two robotic use-cases designed to evaluate its performance in contact-rich scenarios. In our experiments, we deliberately focused on Impedance Control rather than motion generation, since many real-world applications require active compliance, e. g., final assembly tasks where vision is occluded and only tactile sensing is available. The first was a parkour-style task requiring continuous interaction while traversing unknown obstacles. The second was a peg-in-hole insertion task with increasing geometric complexity.
The experiments were conducted on a KUKA LBR iiwa with seven DOFs, using KUKA’s Fast Robot Interface (FRI) for torque control with $5\text{\,}\mathrm{ms}$ sample time. Built-in gravity and Coriolis/centrifugal compensation remained active throughout. Robot kinematics and dynamics were computed via the Exp[licit]-FRI interface [^8] https://GitHub.com/explicit-robotics/Explicit-FRI [35]. A task-space impedance controller was implemented (Section II-B). The external wrenches were acquired with an ATI Gamma force/torque transducer, mounted on the robot flange. The code of the robot controller is publicly available on our public GitHub repository [^9] https://GitHub.com/diffusion-imp-anon/DiffusionBasedImpedanceLearning.
@@
All workpieces were fabricated with a PRUSA i3 MK3 3D printer using PLA filament. The expected manufacturing error of the printer is on the order of $0.1\text{\,}\mathrm{mm}$. All CAD models used in the experiments are available on our GitHub repository. Demonstration videos of both experiments are provided on the project website [^10] https://diffusion-imp-anon.GitHub.io/DiffusionBasedImpedanceLearning/.
All plotted entities are expressed in the spatially-fixed robot base coordinate frame. The colors of the experimental plots are consistent with the colors of the coordinate axes in Figure 9.
Figure 9: Phases of the two experiments and the spatially fixed coordinate frames in which they are expressed. (Left) In the parkour experiments, three obstacles are encountered (numbered 1–3). (Right) During peg insertion, the robot moved from the starting pose to inclined contact with the hole (labeled “Ori”), followed by final insertion (with the start of insertion labeled “Ins”). The colors of the major axes of the coordinate frames are shown in the experimental plots (Figure 10 and 11).
@@ IV-A Parkour
The first experiment consisted of a parkour-style task in which the robot traversed three obstacles while maintaining continuous contact with the table (Figure 9, left). The desired physical interaction behavior was demonstrated by a human operator during data collection (Section III-C), who teleoperated the robot through the parkour task (Figure 6, right).
The nominal ZFT was generated by kinesthetic teaching of a number of key joint poses. A MATLAB script then computed the corresponding end-effector trajectory by smoothly interpolating between these configurations and manually shifting the trajectory along the negative z-direction to ensure constant contact with the table surface (see coordinate frame shown in Figure 9, left). It is important to note that the resulting nominal ZFT passed through the obstacles.
A video of the demonstrations is available on the project website [^11] https://diffusion-imp-anon.GitHub.io/DiffusionBasedImpedanceLearning/ and the MATLAB script used for trajectory generation is publicly available on GitHub [^12] https://GitHub.com/diffusion-imp-anon/DiffusionBasedImpedanceLearning/tree/main/Deployment/TrajectoryGeneration.
@@
To protect the setup, stop conditions of $\lVert\bm{v}\rVert=0.24\ \frac{\text{m}}{\text{s}}$ and $\lVert\bm{f}_{\text{ext}}\rVert=20$ N were defined. The constant-stiffness controller already failed at the first obstacle: the end-effector was stopped, energy accumulated in the virtual spring, and once this energy was released the end-effector accelerated into the velocity stop condition. Forces of up to 15 N were observed. In contrast, with Diffusion-Based Impedance Learning, the robot traversed all obstacles smoothly, i. e., without violating the velocity or force stop conditions. The stiffness adaptation mechanism prevented jamming in directions inconsistent with the ZFT while preserving rigidity along task-relevant axes.
@@
Figure 10 shows the measured external wrenches together with the commanded translational and rotational stiffness values for the Diffusion-Based Impedance Learning. The bottom plots illustrate the relative translational and rotational factors (Equation (34)) and the work done by the external force along the virtual displacement (Equations (25) and (29)).
@@
Figure 10: Measured external wrenches and commanded translational and rotational stiffness during the parkour experiment with stiffness adaptation. The relative translational and rotational factors show their respective contributions to stiffness modulation. The bottom row reports the performed translational and rotational work (Equations (25) and (29)). Shaded gray regions denote the three parkour obstacles. All quantities are expressed in the robot base frame; base-frame color coding and obstacle numbering are shown in Figure 9.
@@
During the entire motion, the stiffness values changed dynamically in response to external wrenches. This continuous adaptation reflected the underlying Diffusion-Based Impedance Learning, which enabled the robot to remain compliant when unexpected contact occurred while maintaining sufficient stiffness to progress through the parkour.
@@
Translational stiffness is adapted asymmetrically depending on the directional contribution of forces. Around 12 s, force peaks of $4\text{\,}\mathrm{N}$ along the $y$-axis and $15\text{\,}\mathrm{N}$ along the $z$-axes arose. Nevertheless, $K_{p,y}$ was reduced more strongly (reduced to near $400\text{\,}\mathrm{N}\mathrm{/}\mathrm{m}$) than $K_{p,z}$ (remaining near $600\text{\,}\mathrm{N}\mathrm{/}\mathrm{m}$) since the relative translational factor along the z-axis was high. A similar effect is visible near $30\text{\,}\mathrm{s}$: while $f_{y}$ reached an absolute of nearly $4\text{\,}\mathrm{N}$ with stiffness near $420\text{\,}\mathrm{N}\mathrm{/}\mathrm{m}$, $f_{z}$ was only $2\text{\,}\mathrm{N}$ yet stiffness dropped to about $200\text{\,}\mathrm{N}\mathrm{/}\mathrm{m}$. These findings confirm that stiffness reduction was not simply proportional to force magnitude but was determined by the alignment with the sZFT, predicted by the Diffusion Model.
@@
Rotational stiffness also adapted directionally, with reductions depending on which axis contributed less to the sZFT. Around $12\text{\,}\mathrm{s}$, moments appeared simultaneously about the $x$- and $y$-axes. Despite higher magnitudes about the $x$-axis, the stiffness $K_{r,y}$ was reduced more strongly, showing that $y$-axis rotations were less important for the intended trajectory. This selective modulation allowed the robot to maintain stability while avoiding unnecessary resistance to contact.
@@ IV-B Peg-In-Hole Insertion
The second experiment was a peg-in-hole insertion task with three different peg types: cylindrical, square, and star pegs. The nominal clearance between each peg and its corresponding hole was $0.14\text{\,}\mathrm{mm}$ for the square peg and $0.20\text{\,}\mathrm{mm}$ for the circular and star pegs. All workpieces were fabricated with the same printer setup as for the parkour experiment. For all trials, task success was defined as completely inserting the peg.
For each peg type, the ZFT was generated from three manually demonstrated robot joint configurations, via kinesthetic teaching: a starting configuration, an initial contact (“touch”) configuration, and a final insertion configuration. These configurations were recorded by physically guiding the robot through the task. A smooth nominal ZFT was then generated using a MATLAB script by computing the corresponding end-effector poses and connecting them with a minimum-jerk interpolation. To induce a controlled pulling force during the final insertion phase, the equilibrium pose for final insertion was deliberately shifted along the negative $z$-direction. After completing the insertion, the robot moved straight upward along the z-axis (see coordinate frame in Figure 9, right). Importantly, during the touch pose the peg’s orientation was not aligned with the edges of the square and star pegs. As a result, the nominal ZFT did not encode any geometric alignment with the peg shape.
A video of the demonstrations is available on the project website [^13] https://diffusion-imp-anon.GitHub.io/DiffusionBasedImpedanceLearning/ and the MATLAB script used for trajectory generation is publicly available on GitHub [^14] https://GitHub.com/diffusion-imp-anon/DiffusionBasedImpedanceLearning/tree/main/Deployment/TrajectoryGeneration.
@@
The three peg types represent increasing geometric complexity. The cylindrical peg is rotation-symmetric, which makes the final orientation about the end-effector $z$-axis irrelevant. In contrast, the square and star pegs require precise rotational alignment with an increasing number of edges, which significantly increases task complexity.
This trend is reflected in the constant-stiffness experiments. The cylindrical peg succeeded in 30/30 trials. The square peg succeeded in 4/30 trials. The star peg failed in all 30 trials.
In practice, more complicated pegs like the square and star peg often require extensive manual tuning of the ZFT, of the impedance parameters, or both. This tuning is not only time-consuming but there is also no standard method that guarantees success. These results show that more complicated assembly processes require more advanced strategies.
In contrast, Diffusion-Based Impedance Learning achieved 30/30 successful insertions for all three peg types. This result was achieved by only demonstrating and interpolating three key poses per trajectory. Hence, no advanced method for trajectory generation was needed. Moreover, the outcome is particularly notable because the training data included only parkour and upper-limb rehabilitation data (Section III-C). No peg-in-hole demonstrations were used for training. Importantly, neither the Diffusion Model nor the impedance controller is provided with any information about the peg or hole geometry, pose, insertion axis, or clearance; all impedance adaptation was driven solely by measured interaction wrenches during the motion. Across all three peg geometries, execution times were nearly identical, as the nominal trajectory was fixed and impedance adaptation operated in parallel.
@@
The videos of the experiments with constant stiffness and Diffusion-Based Impedance Learning can be found on our project website [^15] https://diffusion-imp-anon.GitHub.io/DiffusionBasedImpedanceLearning/.
@@
Figure 11 shows the stiffness adaptation for the star peg. Results for the other two pegs are provided in Appendix B. During all experiments, force and moment thresholds were set to $1$ N and $1$ Nm, respectively.
@@
Figure 11: Measured external wrenches and commanded translational and rotational stiffness during the peg-in-hole star insertion experiment with stiffness adaptation. The relative translational and rotational factors show their respective contributions to stiffness modulation. The bottom row reports the performed translational and rotational work (Equations (25) and (29)). The shaded gray region denotes the insertion phase. All quantities are expressed in the robot base frame; base-frame color coding is shown in Figure 9.
@@
As illustrated in Figure 11, stiffness values were reduced in response to external wrenches, with the amount of reduction governed by the relative translational and rotational factors. For most of the experiment, the relative translation factor along the $z$ axis remained higher than along the $y$ axis. For instance, at around 11 $1\text{\,}\mathrm{s}$, small forces of $0.5\text{\,}\mathrm{N}$ appeared along the $y$ axis. Nevertheless, the stiffness $K_{p,y}$ was reduced to $600\text{\,}\mathrm{N}\mathrm{/}\mathrm{m}$. During retraction of the peg at around $16\text{\,}\mathrm{s}$, it can be seen that the relative translation factor along the $z$ axis dropped and consequently the stiffness $K_{p,z}$ was reduced to $500\text{\,}\mathrm{N}\mathrm{/}\mathrm{m}$. In contrast, the relative translation factor along the $x$ axis increased and consequently the stiffness $K_{p,x}$ was increased to $700\text{\,}\mathrm{N}\mathrm{/}\mathrm{m}$.
[…]
IV-C Ablation Study
In addition to the two main controllers, we performed an ablation study to isolate the contributions of the reconstructed sZFT (Section III-A) and the directional impedance adaptation (Section III-E).
Importantly, the nominal ZFT was kept fixed throughout the ablation study. It was generated from simple kinesthetic demonstrations and was not optimized for the two tasks and pegs. Videos illustrating the kinesthetic teaching procedure for the parkour and peg-in-hole tasks are provided on the project website.
Specifically, we evaluated the following variants using the baseline impedance parameters ($800~\text{N/m}$ for translation and $150~\text{Nm/rad}$ for rotation):
  1. 1.
    Uniform stiffness adaptation without directional adaptation, using the reconstructed sZFT
  2. 2.
    Directional stiffness adaptation computed from the nominal ZFT
@@ IV-C Ablation Study
Uniform stiffness adaptation without directional modulation failed in both experimental scenarios. In the parkour task, the robot was blocked at the first obstacles, while none of the peg-in-hole insertions succeeded. Directional adaptation derived from the nominal ZFT also failed in contact-rich settings. In the parkour task, the robot exceeded safety limits and stopped. In the peg-in-hole experiments, this variant achieved $30/30$ success only for the cylindrical peg, but failed for both the square and star pegs.
These results confirm that directional adaptation alone is insufficient if it is derived from the nominal ZFT. Instead, directional factors must be computed relative to the reconstructed sZFT for impedance adaptation to succeed.
To further evaluate how different uniform and non-uniform fixed impedance parameters influence task success in the peg-in-hole experiment, we additionally evaluated the following variants:
  1. a)
    Uniform high impedance: $800~\text{N/m}$ (translation), $150~\text{Nm/rad}$ (rotation)
  2. b)
    Uniform medium impedance: $400~\text{N/m}$ (translation), $50~\text{Nm/rad}$ (rotation)
  3. c)
    Uniform low impedance: $200~\text{N/m}$ (translation), $7~\text{Nm/rad}$ (rotation)
  4. d)
    Best-found, manually tuned non-uniform impedance: translation along $z$: $800~\text{N/m}$; translation along $x$ and $y$: $400~\text{N/m}$; rotation about $z$: $5~\text{Nm/rad}$; rotation about $x$ and $y$: $100~\text{Nm/rad}$
For a) (uniform high impedance), the cylindrical peg succeeded in $30/30$ trials, the square peg in $4/30$ trials, and the star peg failed in all trials. For b) (uniform medium impedance), yielded $30/30$ success for the cylindrical peg, $13/30$ success for the square peg, and $5/30$ success for the star peg. For c) (uniform low impedance), it failed for all peg types, as the wrench required to orient the peg with the hole was insufficient.
For d) (manually tuned non-uniform impedance), the cylindrical peg succeeded in $30/30$ trails, the square peg in $26/30$ trials, and the star peg in $21/30$ trials. These parameters were identified by an experienced impedance control programmer after testing 15 different parameter combinations. In contrast, inexperienced users would likely struggle to identify a single parameter set that performs robustly across all peg geometries.
Videos of all ablation experiments for the star peg example are provided on our project website.
A summary of the ablation study for the peg-in-hole experiment can be found in Table III.
TABLE III: Peg-in-hole success rates for the ablation study. All results are reported over 30 trials.
Controller Variant Cylindrical Peg Square Peg Star Peg
Time-varying Impedance
[…]
Diffusion Model w/o directional adaptation 0% 0% 0%
Directional adaptation w/o DM (nominal ZFT) 100% 0% 0%
Fixed Impedance
Uniform high (baseline) 100% 13% 0%
Uniform medium 100% 43% 17%
Uniform low 0% 0% 0%
Best-found non-uniform 100% 87% 70%